The velocity of sound in a gas of pressure p and density roll is written as v=kp^xp^y find x and t and state the relationship

1
Expert's answer
2022-01-11T04:13:16-0500

The velocity of a sound in gas is related to the pressure "p" and density "\rho" by the formula

"v=kp^x\rho^y"

We use the dimensional analysis to find "x" and "y".

"[v]=LT^{-1}\\
[p]=MT^{-2}L^{-1}\\
[\rho]=ML^{-3}"

We get

"LT^{-1}=(MT^{-2}L^{-1})^x (ML^{-3})^y"

or

"-x-3y=1,\\
x+y=0,\\
-2x=-1."

Roots:


"x=1/2,\quad y=-1/2"

So, the velocity of a sound in gas is given by

"v=kp^{1/2}\rho^{-1/2}=k\sqrt{p/\rho}"